### v1

Abstract.
We prove rapid mixing for almost all random walks generated by $m$ translations on an arbitrary nilmanifold under mild assumptions on the size of $m$. For several classical classes of nilmanifolds, we show $m=2$ suffices. This provides a partial answer to the question raised in [6] about the prevalence of rapid mixing for random walks on homogeneous spaces.
1. Introduction
Let $G$ be a simply connected nilpotent Lie group and $\Gamma$ be a co-compact lattice so that $M=G/\Gamma$ is a nilmanifold equipped with the Haar measure $\mu$. In this paper, we study random walks by a finite set of translations on $M$. These random walks will be defined by an $m\in{\mathbb{N}},\ x\in M$, a set $F:=\{g_{1},\ldots,g_{m}\}\subset G$, and an associated probability vector $\vec{p}=(p_{1},p_{2},\dots,p_{m})$, i.e. $p_{i}>0$ and $\sum p_{i}=1$. The random walk is then a Markov chain where $x_{n}=g_{k}x_{n-1}$ with the probability $p_{k}$. Associated to the set $F$, we have an operator ${\mathcal{L}}:C^{r}(M)\rightarrow C^{r}(M)$ defined by
${\mathcal{L}}(A)(x):={\mathbb{E}}(A(x_{1})|x_{0}=x)=\sum_{j=1}^{m}p_{j}A(g_{j}\cdot x),\ A\in C^{r}(M).$
It follows that for any $N>1$, $({\mathcal{L}}^{N}A)(x)={\mathbb{E}}_{x}(A(x_{N})).$ Since a random walk is exactly defined by its generators, we can view the product $\mu^{m}$ as a measure on the space of all random walks generated by $m$ translations. When we speak of the measure of a set of walks, we mean it with respect to this measure.
Given observables $A,B\in C^{r}(M)$, the correlation of $A$ and $B$ after time $N$ is given by
$\overline{\rho}_{A,B}(N)=\int({\mathcal{L}}^{N}A)(x)B(x)d\mu(x)-\int A(x)d\mu(x)\int B(x)d\mu(x).$
Definition 1.1.
A random walk is said to be rapid mixing if given $q\in{\mathbb{N}}$, there are constants $C,r>0$ such that for any $A,B\in C^{r}(M)$ and $N\in{\mathbb{N}}$,
$|\overline{\rho}_{A,B}(N)|\leq C\left\|A\right\|_{C^{r}(M)}\left\|B\right\|_{C^{r}(M)}N^{-q}.$
Below, in order to simplify the formulas, we only consider correlations of zero mean functions. This is sufficient since every function can be decomposed as a sum of a zero mean function and a constant (the mean).
Mixing plays a key role in the study of statistical properties of dynamical systems. Many classical systems are exponentially mixing (see e.g. [7, Appendix A]). The random walks we consider do not exhibit exponential mixing, so the best estimate we can hope for is that the mixing of $C^{r}$ observables occurs at a $O(N^{-q(r)})$ rate with $q(r)\to\infty$ as $r\to\infty$. This is exactly the definition of rapid mixing that we have provided. Our main result shows that most sufficiently rich walks are indeed rapid mixing. Rapid mixing is also sufficient to establish several key statistical properties, including the Central Limit Theorem, see Appendix B.
Nilmanifolds support a rich and well-studied variety of homogeneous dynamical systems. These systems, in addition to being of purely dynamical interest, are also of interest to the broader mathematical community as their dynamical properties can have important consequences in other fields, particularly number theory [13, 10] and combinatorics [14]. In part due to the complexity of their algebraic structures, dynamics on higher-dimensional nilmanifolds are not yet completely understood. In this paper, we prove the rapid mixing of almost all random walks on nilmanifolds generated by a sufficiently large (finite) number of group elements. In particular, we will define a technical algebraic condition called $m$-greatness and show that almost any random walk on an $m$-great nilmanifold supported on $m$ translations is rapid mixing. We then show that every step-$s$ nilmanifold is $s$-great. This leads to the following theorem which applies to all nilmanifolds.
Theorem 1.2.
For any step-$s$ nilmanifold $M=G/\Gamma$, there is a constant $N_{G}\leq s$ such that for $m\geq N_{G}$, almost every random walk generated by $m$ translations is rapid mixing.
The primary weakness of this theorem is that it sometimes requires more than two generators to guarantee that rapid mixing random walks are of full measure. Within certain special classes of nilmanifolds, we can overcome this deficiency. Among these are the quasi-abelian nilmanifolds, which have previously been studied in the context of parabolic flow dynamics [19, 12]. We also give special attention to triangular nilmanifolds and step-3 nilmanifolds, and ultimately show the following theorem.
Theorem 1.3.
If $G$ is quasi-abelian, triangular, or step-3 or lower, then $N_{G}=2$.
As we shall see, the argument requires certain polynomials arising from the Lie algebra structure to be linearly independent.
A rapid mixing random walk on $M$ also satisfies the Central Limit Theorem (see Appendix B). Several results on central and local limit theorems on nilpotent Lie groups were recently obtained in [15, 2, 3, 5].
We now describe the structure of the paper. Section 2 contains preliminaries on nilpotent Lie groups, Lie algebras, and nilmanifolds. In Section 3, we define $m$-greatness and show that it implies rapid mixing for a full measure set of $m$-tuples modulo a lemma that connects $m$-greatness to Diophantine properties. In Section 4, we show the lemma by constructing words that act Diophantinely on appropriate tori. In Section 5, we establish that the groups listed in Theorem 1.3 are 22-great implying Theorem 1.3. We also show that any step $s$-nilmanifold is $s$-great, proving Theorem 1.2. Finally, we provide an example of a Lie algebra that is not $2$-great, showing that our technique is not sufficient to show the suspected optimal result that $N_{G}\!\!=\!\!2$ for all nilpotent groups.
Acknowledgement
The authors greatly appreciate Adam Kanigowski for fruitful discussions and several suggestions for improving the quality of the initial draft. They also thank Bassam Fayad and Giovanni Forni for their several comments on the draft. M.K. thanks Xuesen Na for his assistance in setting up a computer program for testing calculations.
M.K. is grateful to the Institut de Mathématiques de Jussieu-Paris Rive Gauche and the Center of Excellence “Dynamics, Mathematical Analysis and Artificial Intelligence” at Nicolaus Copernicus University in Toruń for their hospitality during his visits.
D.D. was partially supported by the NSF grant DMS-2246983.
S.D. was partially supported by the NSF grant DMS-2101464.
M.K. was partially supported by UniCredit Bank R&D group through the ‘Dynamics and Information Theory Institute’ at the Scuola Normale Superiore, foundations for the Royal Swedish Academy of Sciences, and Carl Trygger’s Foundation for Scientific Research.
2. Preliminaries on nilmanifolds
We provide a background of nilpotent Lie groups, Lie algebras and nilmanifolds. The material of this subsection is taken from [4, 18], and we also refer to [1] for additional information about general nilmanifolds.
2.1. Nilpotent Lie groups and Lie algebras
A real Lie algebra $\mathfrak{g}$ is called nilpotent if the lower (descending) central series of $\mathfrak{g}$ terminates i.e. the sequence defined by
(2.1) $\mathfrak{g}=\mathfrak{g}^{(0)}\supset\mathfrak{g}^{(1)}=[\mathfrak{g},\mathfrak{g}]\supset\cdots\supset\mathfrak{g}^{(j)}=[\mathfrak{g}^{(j-1)},\mathfrak{g}]\supset\cdots,$
where $\mathfrak{h},\mathfrak{g}]=\{[X,Y]:X\in\mathfrak{h},Y\in\mathfrak{g}\}$, eventually has $\mathfrak{g}^{(s)}=0$ for some $s$. The step of $\mathfrak{g}$ is the minimal number $s$ that satisfies $\mathfrak{g}^{(s)}=0$.
The lower central series of a Lie group $G$ is defined by $G^{(0)}=G$ and $G^{(j+1)}=[G^{(j)},G]$ where $\cdot$$\cdot$] is the commutator bracket. A connected and simply connected Lie group $G$ is called nilpotent if $G^{(s)}$ is equal to the trivial group for some $s$. A Lie group $G$ is nilpotent if and only if its associated Lie algebra $\mathfrak{g}$ is nilpotent. In fact, the lower central series of $G$ and $\mathfrak{g}$ are connected, as $G^{(j)}=\exp(\mathfrak{g}^{(j)})$.
Proposition 2.1.
[4] If $G$ is a connected and simply connected nilpotent Lie group, then exponential map $\exp:\mathfrak{g}\rightarrow G$ is a diffeomorphism.
The product operation on $G$ satisfies the Baker-Campbell-Haussdorff (BCH) formula
(2.2) $\exp(X)\exp(Y)=\exp\left(X+Y+\frac{1}{2}[X,Y]+\sum_{\alpha}c_{\alpha}X_{\alpha}\right),$
where $\alpha$ is a finite (for nilpotent groups) set of labels, $c_{\alpha}$ are real constants, and $X_{\alpha}$ are iterated Lie brackets of $X$ and $Y$ (see [9]).
It will be convenient to denote $n_{j}=\dim(\mathfrak{g}^{(j)})-\dim(\mathfrak{g}^{(j+1)})$ so that $n_{j}$ is the dimension of the quotient algebra $\mathfrak{g}^{(j)}/\mathfrak{g}^{(j+1)}$ (or corresponding quotient group $G^{(j)}/G^{(j+1)}$). If $X,Y\in\mathfrak{g}$ satisfy $X-Y\in\mathfrak{g}^{(j)}$ we write that $X=Y\mod\mathfrak{g}^{(j)}$.
Definition 2.2 (Malcev basis).
A Malcev basis for $\mathfrak{g}$ through the descending central series $\mathfrak{g}^{(j)}$ is a basis $X_{1}^{(0)},\cdots X_{n_{0}}^{(0)},$ $\cdots,X_{1}^{(s-1)},\cdots,X_{n_{s}}^{(s-1)}$ of $\mathfrak{g}$ satisfying the following:
(1) if we set $E^{j}=\{X_{1}^{(j)},\cdots,X_{n_{j}}^{(j)}\}$, the elements of the set $E^{j}\cup E^{j+1}\cup\cdots\cup E^{s}$ form a basis of $\mathfrak{g}^{(j)}$;
(2) if we drop the first $l$ elements, the remaining elements span an ideal (of codimension $l$) of $\mathfrak{g}$.
If $\Gamma=\displaystyle\left\{\exp\left(\sum_{j,k}m_{j,k}X_{j}^{(k)}\right)\right\}_{m_{j,k}\in{\mathbb{Z}}}$, then we say that the basis is strongly based at $\Gamma$. The Lie algebra $\mathfrak{g}$ of any nilmanifold $G/\Gamma$ can be equipped with a Malcev basis strongly based at $\Gamma$. Moving forward, we will always use this basis when writing Lie algebra elements in coordinates. For convenience, we will denote by $X_{i}=X_{i}^{(0)}$.
2.2. Nilmanifolds and Fibration
A compact nilmanifold is a quotient $M:=G/\Gamma$ where $G$ is a nilpotent group and $\Gamma$ is a (co-compact) lattice of $G$. The lattice $\Gamma$ exists if and only if $G$ admits rational structural constants. We will consider the left action of $G$ by translations on $M$. More precisely, for $g,h\in G$, set $g(h\Gamma)=(gh)\Gamma.$ Every nilmanifold is a fiber bundle over a torus. The abelianization $G^{ab}=G/[G,G]$ is abelian, connected and simply connected, hence isomorphic to ${\mathbb{R}}^{n}$. Thus, there is a natural projection
(2.3) $p:G/\Gamma\rightarrow G^{ab}/\Gamma^{ab}\simeq{\mathbb{T}}^{n}.$
For all $k\in{\mathbb{N}}_{0}$, the group $G^{(k+1)}$ is a closed normal subgroup of $G$, we have natural epimorphisms $\pi^{(k)}:G\rightarrow G/G^{(k+1)}$. Then, the group $G^{(k+1)}\cap\Gamma$ is a lattice of $G^{(k+1)}$. Moreover, $\Gamma_{k}:=\pi^{(k)}(\Gamma)$ is a lattice in $N_{k}:=G/G^{(k+1)}$ and
$M^{(k)}:=G/G^{(k+1)}\Gamma=N_{k}/\Gamma_{k}$
is a nilmanifold. It follows that $\pi^{(k)}:M=G/\Gamma\rightarrow M^{(k)}=G/G^{(k+1)}\Gamma$ is a fibration whose fibers are the orbits of $G^{(k+1)}$ on $G/\Gamma$, homeomorphic to the nilmanifolds $G^{(k+1)}/(G^{(k+1)}\cap\Gamma)$. We can also define $M_{p}:=G^{(p)}/G^{(p+1)}\Gamma_{p}\simeq{\mathbb{T}}^{n_{p}}$.
2.3. Main examples
In this article, we will pay special attention to two classes of Lie groups. The first class was first in [20], (see also [10] for general introduction).
Definition 2.3.
A Lie algebra $\mathfrak{g}$ is called quasi-abelian if it is not abelian and has an abelian subalgebra of codimension 1.
It follows from Definition 2.3 that any quasi-ableian Lie algebra $\mathfrak{g}$ has a basis $(X,Y_{i,j})_{(i,j)\in J}$, satisfying the commutation relations
$X,Y_{i,j}]=Y_{i+1,j},\ (i,j)\in J,$
and all other commutation relations are trivial (see [11]). We also remark that the class of quasi-abelian Lie algebra contains the class of filiform Lie algebras (see [20]), so our results also hold in the filiform case.
The next example is not quasi-abelian, but it has a tractable structure.
Definition 2.4.
A Lie algebra is triangular if it isomorphic to $\mathfrak{t}_{s}$–the Lie algebra of strictly upper triangular $(s+1)\!\times\!(s+1)$ matrices with the standard bracket for some $s$.
It follows from the definition that $\mathfrak{t}_{s}$ is $\frac{1}{2}s(s+1)$-dimensional. Letting $E_{ij}$ represent the $(s+1)\times(s+1)$ matrix with a 1 in position $(i,j)$ and zeroes elsewhere, we see that $\{E_{ij}:j>i\}$ forms a basis for $\mathfrak{t}_{s}$. The relationship among these basis elements are given by
$E_{ij},E_{i^{\prime}j^{\prime}}]=\delta_{ji^{\prime}}E_{ij^{\prime}}-\delta_{j^{\prime}i}E_{i^{\prime}j}.$
We refer to [4, 16] for additional information on triangular algebras.
2.4. Harmonic Analysis on Nilmanifolds
In this section, we wish to describe the structure of $L^{2}(M)$ for a nilmanifold $M$. The first step is to describe the characters of $M$, denoted by $\widehat{M}$. For any Lie algebra $\mathfrak{a}$ let $\mathfrak{a}^{*}$ represent its dual as a vector space. Since we have a basis for $\mathfrak{g}$ (the Malcev basis), the coordinate functionals $X_{1}^{(1)*},\ldots,X_{n_{s-1}}^{(s-1)*}$ are well-defined and form a basis for $\mathfrak{g}^{*}$. Define $\pi_{p}:\mathfrak{g}\rightarrow{\mathbb{R}}^{n_{p}}$ by $\pi_{p}(V)=(X_{1}^{(p)*}(V),\ldots,X_{n_{p}}^{(p)*}(V))$. Also, let $E_{p}^{*}=\{X_{1}^{(p)*},\ldots,X_{n_{p}}^{(p)*}\}$, and define $\Lambda_{p}$ to be the set of integer linear combinations of elements of $E_{p}^{*}$. We let $\Lambda=\bigoplus_{i=0}^{s-1}\Lambda_{i}.$ Precomposing an element of $\Lambda$ with $\log$ and postcomposing by the complex exponential $e(x)=e^{2\pi ix}$ gives a character on $M$. In fact, $\widehat{M}$ is exactly equal to $e\circ\Lambda\circ\log$. When $\lambda\in\Lambda$, we will let $\chi_{\lambda}$ denote the character defined by $\chi_{\lambda}(g)=e^{2\pi i\lambda(\log g)}$. We define a a norm on $\widehat{M}$ by $\|\chi_{\lambda}\|=\|\lambda^{*}\|$ where $\lambda^{*}$ is the element of $\mathfrak{g}$ dual to $\lambda$. Finally, for any characters $\chi\in\widehat{M}$, we let
$H_{\chi}=\{\varphi\in L^{2}(M):\varphi(gx)=\chi(g)\varphi(x)\}.$
We have that
$L^{2}(M)=\bigoplus_{\chi\in\widehat{M}}H_{\chi}.$
Moreover, when $\varphi\in C^{r}(G/\Gamma)$, if we write $\varphi=\sum_{\chi\in\widehat{M}}\varphi_{\chi}$ with $\varphi_{\chi}\in H_{\chi},$ then the weight functions $\varphi_{\chi}$ satisfy $\|\varphi_{\chi}\|_{C^{0}}\leq\frac{\|\varphi\|_{C^{r}}}{\|\chi\|^{r}}.$
2.5. Diophantine Conditions
Given a vector $v\in{\mathbb{R}}^{d}$, we say that $v\in DC(\gamma,\tau)$ if for all $n\in{\mathbb{Z}}^{d}$ and $m\in{\mathbb{Z}}$
$|n\cdot v-m|\geq\frac{\gamma}{|n|^{\tau}}.$
In [17], Kleinbock and Margulis show that within submanifolds of ${\mathbb{R}}^{n}$ satisfying a linear independence condition almost every point is in $DC(\gamma,\tau)$ for some $\gamma$ and $\tau.$ In particular, they show the following which is listed as Conjecture $H_{1}$ in their paper.
Theorem 2.5.
[17] Let $f_{1},\ldots,f_{n}$ along with the constant function 1 form a linearly independent set of analytic functions from ${\mathbb{R}}^{m}$ to $\mathbb{R}$}. For almost every $x\in{\mathbb{R}}^{m}$, there exist $\gamma,\tau>0$ such that $(f_{1}(x),\ldots,f_{n}(x))\in DC(\gamma,\tau)$.
This result will provide the bridge from the Lie algebra structure to the Diophantine estimates needed to show mixing. Since the functions we will consider will always be homogeneous polynomials with positive degree, linear independence of $f_{1},...,f_{n}$ is the only condition we will need to check in order to apply this theorem in practice.
3. $m$-greatness and mixing
In this section, we will first define what it means for a nilmanifold to be $m$-great and then show that on these manifolds almost any $m$-tuple generates a rapidly mixing walk.
3.1. Linear independence of polynomials
Let $\mathcal{H}(t,\alpha)=(\mathcal{H}_{1}(t,\alpha),\dots,\mathcal{H}_{\ell}(t,\alpha))$ be a polynomial map of ${\mathbb{R}}^{a}\times{\mathbb{R}}^{b}\to{\mathbb{R}}^{\ell}.$ Let
$\mathcal{D}=\{\bar{t}\in{\mathbb{R}}^{a}:\mathcal{H}_{1}(\bar{t},\alpha),\dots\mathcal{H}_{\ell}(\bar{t},\alpha)\text{ are linearly dependent}\}.$
Expanding $\mathcal{H}_{j}(\bar{t},\alpha)=\sum_{{\mathbf{m}}}c_{j,{\mathbf{m}}}(\bar{t})\alpha^{{\mathbf{m}}}$ we see that the above conditions amounts to vanishing of certain minors of the matrix $c_{j,{\mathbf{m}}}$ whence $\mathcal{D}$ is an algebraic subvariety of ${\mathbb{R}}^{a}.$ We say that $\mathcal{H}$ is non-degenerate if $\mathcal{D}\neq{\mathbb{R}}^{a}.$
Now we describe special polynomial mapping associated to a Lie algebra. Fix integers $m$ and $p.$ Consider $m$ vectors $V_{i}=\sum_{j=1}^{n}\alpha_{ij}X_{j},$ $i\in\{1,\dots m\}$ and let
$\mathcal{H}_{m,p}(t,\alpha)=[\ldots\bf{[t_{0}V,t_{1}V],\ldots],t_{p}V]}+\mathfrak{g}^{(p+1)}\in\mathfrak{g}^{(p)}/\mathfrak{g}^{(p+1)}$
where ${\mathbf{t}}_{q}=(t_{q1},\dots t_{qm})$ and ${\mathbf{t}}\textbf{V}=\sum_{i=1}^{m}t_{i}V_{i}.$ Note that we have
(3.1) $\mathcal{H}_{m,p}({\mathbf{k}},\alpha)=\sum_{i_{j}\in\{1,\ldots,m\}}k_{0i_{0}}\ldots k_{pi_{p}}M_{i_{0}i_{1},\ldots,i_{p}}+\mathfrak{g}^{(p+1)}$
where $M_{i_{0}i_{1},\ldots,i_{p}}=[\ldots[V_{i_{0}},V_{i_{1}}],\ldots],V_{i_{p}}].$ In coordinates, we see that
$\mathcal{H}_{m,p}({\mathbf{k}},\alpha)=\sum_{i=1}^{n_{p}}P_{i}(k,\alpha)X_{i}^{(p)}+\mathfrak{g}^{(p+1)}$
where each $P_{i}({\mathbf{k}},\alpha)$ is a polynomial in the coordinates of $\mathbf{k}$} and $\alpha_{ij}$. From now on we will suppress the $\mathfrak{g}^{(p+1)}$ that ought to appear whenever we discuss $\mathcal{H}_{2,p}$, but we agree that $\mathcal{H}_{2,p}$ is always defined modulo $\mathfrak{g}^{(p+1)}$. We now provide a brief illustrative example.
Example 3.1.
Let $\mathfrak{g}$ be the step-3 Lie algebra of dimension 5 with the following commutation relations
$[X_{1},X_{2}]=Y,\quad[Y,X_{1}]=Z_{1},\quad[Y,X_{2}]=Z_{2}.$
with all other brackets being $0.$ Let $V_{1}=\alpha_{11}X_{1}+\alpha_{12}X_{2}+\dots$, $V_{2}=\alpha_{21}X_{1}+\alpha_{22}X_{2}+\dots.$ Then $[[V_{1},V_{2}],V_{1}]=(\alpha_{11}^{2}\alpha_{22}-\alpha_{11}\alpha_{12}\alpha_{21})Z_{1}+(\alpha_{11}\alpha_{12}\alpha_{21}-\alpha_{12}^{2}\alpha_{21})Z_{2}\mod\mathfrak{g}_{3}.$ Therefore $M_{121}=\begin{bmatrix}\alpha_{11}^{2}\alpha_{22}-\alpha_{11}\alpha_{12}\alpha_{21}\\ \alpha_{11}\alpha_{12}\alpha_{21}-\alpha_{12}^{2}\alpha_{21}\end{bmatrix}$.
Suppose now we wish to compute the $\mathcal{H}_{2,2}$ for this $\mathfrak{g}$. To do so, we can proceed in two different ways. We could compute each $M_{i_{1}i_{2}i_{3}}$ and then write the sum as in (3.1). Alternatively, we can explicitly compute $\bf{[k_{0}V,k_{1}V],k_{2}V]}$ and we shall see how the $M_{i_{1}i_{2}i_{3}}$ appear. We will take this second approach.
$[[k_{01}V_{1}+k_{02}V_{2}$ $,k_{11}V_{1}+k_{12}V_{2}],k_{21}V_{1}+k_{22}V_{2}]$
$=[(k_{01}k_{12}$ $(\alpha_{11}\alpha_{22}-\alpha_{12}\alpha_{21})-k_{02}k_{11}(\alpha_{11}\alpha_{22}-\alpha_{12}\alpha_{21}))Y,k_{21}V_{1}+k_{22}V_{2}]$
$=k_{01}k_{12}k_{21}\big((\alpha_{11}^{2}\alpha_{22}-\alpha_{11}\alpha_{12}\alpha_{21})Z_{1}+(\alpha_{11}\alpha_{12}\alpha_{22}-\alpha_{12}^{2}\alpha_{21})Z_{2}\big)$
$+k_{01}k_{12}k_{22}\big((\alpha_{11}\alpha_{21}\alpha_{22}-\alpha_{12}\alpha_{21}^{2})Z_{1}+(\alpha_{11}\alpha_{22}^{2}-\alpha_{12}\alpha_{21}\alpha_{22})Z_{2}\big)$
$-k_{02}k_{11}k_{21}\big((\alpha_{11}^{2}\alpha_{22}-\alpha_{11}\alpha_{12}\alpha_{21})Z_{1}+(\alpha_{11}\alpha_{12}\alpha_{22}-\alpha_{12}^{2}\alpha_{21})Z_{2}\big)$
$-k_{02}k_{11}k_{22}\big((\alpha_{11}\alpha_{21}\alpha_{22}-\alpha_{12}\alpha_{21}^{2})Z_{1}+(\alpha_{11}\alpha_{22}^{2}-\alpha_{12}\alpha_{21}\alpha_{22})Z_{2}\big).$
We see that $M_{i_{1}i_{2}i_{3}}$ is exactly the coefficient of $k_{0i_{1}}k_{1i_{2}}k_{2i_{3}}$, and our mapping can be written as
$k_{01}k_{12}k_{21}M_{121}+k_{01}k_{12}k_{22}M_{122}+k_{02}k_{11}k_{21}M_{211}+k_{02}k_{11}k_{22}M_{212}.$
In vector form, $\mathcal{H}_{2,2}$ takes form
$k_{01}k_{12}k_{21}\begin{bmatrix}\alpha_{11}^{2}\alpha_{22}-\alpha_{11}\alpha_{12}\alpha_{21}\\ \alpha_{11}\alpha_{12}\alpha_{21}-\alpha_{12}^{2}\alpha_{21}\end{bmatrix}+k_{01}k_{12}k_{22}\begin{bmatrix}\alpha_{11}\alpha_{21}\alpha_{22}-\alpha_{12}\alpha_{21}^{2}\\ \alpha_{11}\alpha_{22}^{2}-\alpha_{12}\alpha_{21}\alpha_{22}\end{bmatrix}$
$-k_{02}k_{11}k_{21}\begin{bmatrix}\alpha_{11}^{2}\alpha_{22}-\alpha_{11}\alpha_{12}\alpha_{21}\\ \alpha_{11}\alpha_{12}\alpha_{21}-\alpha_{12}^{2}\alpha_{21}\end{bmatrix}-k_{02}k_{11}k_{22}\begin{bmatrix}\alpha_{11}\alpha_{21}\alpha_{22}-\alpha_{12}\alpha_{21}^{2}\\ \alpha_{11}\alpha_{22}^{2}-\alpha_{12}\alpha_{21}\alpha_{22}\end{bmatrix}$.
We clearly see that $\mathcal{H}_{2,2}$ is non degenerate since even just setting $k_{01}=k_{12}=k_{21}=1$ and $k_{02}=k_{11}=k_{22}=0$, we get that at that value of $\mathbf{k}$} $\mathcal{H}_{2,2}$ evaluates to $M_{121}$, which has linearly independent rows.
3.2. $m$-great Lie algebras and mixing
Now, we will define the key technical property which guarantees that the set of rapid mixing $m$-tuples is of full measure.
Definition 3.2.
If for every $i\in[1,s-1]$, the $\mathcal{H}_{m,i}$ is non degenerate we say that $\mathfrak{g}$ is mm–great.
When a Lie algebra is not mm–great, we say it is mm–bad. We will call a Lie group or a nilmanifold $m$-great if its associated Lie algebra is $m$-great. We will show that $m$-greatness implies that the walk generated almost every $m$-tuple has a Diophantine property in the following sense.
Definition 3.3.
Let $M=G/\Gamma$ be a nilmanifold and $(g_{1},\ldots g_{m})$ be an $m$-tuple.
  • A pair of words $W_{1},W_{2}\in\langle g_{1},\ldots,g_{m}\rangle$ is said to be nice on $G^{(p)}$ if they have equal length and their difference $h:=W_{2}W_{1}^{-1}$ satisfies $h\in G^{(p)}$, and there exist $\gamma,\tau>0$ such that $\pi_{p}(\log(h))\in DC(\gamma,\tau)$.
  • An $m$-tuple is said to act Diophantinely on $M_{p}$ if there exists a pair of words that is nice on $G^{(p)}$.
  • We say that an $m$-tuple acts Diophantinely at all levels if it acts Diophantinely on $M_{p}$ for all $0\leq p<s.$
Let $E_{m}$ denote the set of $m$-tuples which act Diophantinely at all levels on a nilmanifold $M$, and recall that $\mu^{m}$ is the measure on the space of $m$-tuples of elements of $G$ given by the $m$-fold product of the Haar measure on $G$ with itself.
Lemma 3.4.
If $M=G/\Gamma$ is $m$-great, then $E_{m}$ is of full measure with respect to $\mu^{m}$.
We postpone the proof to Section 4. We now show that the random walk generated by any $m$-tuple that act Diophantinely on all levels is rapid mixing. The key estimate to show appears in the following theorem:
Theorem 3.5.
If a random walk is generated by an $m$-tuple that acts Diophantinely at all levels, then there exist constants $C,a,b>0$ such that for $A\in C^{r}(M)$ with zero mean,
(3.2) $\|{\mathcal{L}}^{N}A\|_{C^{0}}\leq\frac{K\|A\|_{C^{r}}}{N^{ar-b}}.$
Whenever (3.2) holds, it implies rapid mixing of the associated random walk since
(3.3) $|\overline{\rho}_{A,B}(N)|\leq\|{\mathcal{L}}^{N}A\|_{C^{0}}\|B\|_{C^{0}}\leq C\left\|A\right\|_{C^{r}}\left\|B\right\|_{C^{0}}N^{-(ar-b)}.$
To prove this Theorem 3.5, we will show that for any nontrivial character $\chi$, $\mathcal{L}$} has spectral gap on $H_{\chi}$, and, moreover, that the size of the spectral gap has a polynomial lower bound as $\chi$ grows in norm. We will also use the fact that the size of a $C^{r}$ function’s projection into $H_{\chi}$ is polynomially bounded in the norm of $\chi$. Together, these estimates will allow us to prove Theorem 3.5. We now show the spectral gap.
Lemma 3.6.
Assume that $(g_{1},...,g_{m})$ acts Diophantinely at all levels. There exist constants $C_{1},C_{2},\ell,\tau>0$ such that if $\chi$ is a nontrivial character on $M$ and $A\in H_{\chi}$, then
$\|{\mathcal{L}}^{N}A\|_{C^{0}}\leq C_{1}\left(1-C_{2}\left\|\chi\right\|^{-2\tau}\right)^{N/\ell}\|A\|_{C^{0}}.$
Proof.
Let $\lambda=\lambda_{0}\oplus\ldots\oplus\lambda_{s-1}\in\Lambda$ be such that $\chi=\chi_{\lambda}$. Let $i$ be the largest value such that $\lambda_{i}\neq 0$. Let $W_{1}$ and $W_{2}$ be nice on $G^{(i)}$. Let $\ell_{i}$ be the common length of $W_{1}$ and $W_{2}$, and let $\gamma_{i},\tau_{i}$ be the associated Diophantine constants. We note that each of these constants depend only on $i$ and not on the particular character $\chi$. Let ${\mathbf{p}}=\min(p(W_{1}),p(W_{2})).$ Suppose without loss of generality that ${\mathbf{p}}=p(W_{1})$. Then setting $h=W_{2}W_{1}^{-1}$
${\mathcal{L}}^{\ell_{i}}(A)(x)={\mathbf{p}}\left[A(W_{1}x)+A(hW_{1}x)\right]+\left[(p(W_{2})-p(W_{1}))A(W_{2}x)+\sum_{j\geq 3}p(W_{j})A(W_{j}x)\right].$
By Lemma A.1 from the appendix, we have that the first term satisfies
${\mathbf{p}}|A(W_{1}x)+A(hW_{1}x)|={\mathbf{p}}|1+\chi(h)||A(W_{1}x)|\leq 2{\mathbf{p}}(1-{\bar{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}},$
while the second term is at most $(1-2{\mathbf{p}})\|A\|_{C^{0}}.$ Combining these estimates, we see that on $H_{\chi}$,
$\|{\mathcal{L}}^{\ell_{i}}A\|_{C^{0}}\!\!\leq\!\!(1\!-\!{\hat{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}}.$
Fixing ${\hat{C}}=\min\{{\hat{C}}_{i}\}$, $\ell=\max\{\ell_{i}\}$, and $\tau=\max\{\tau_{i}\}$, we see that
$\|{\mathcal{L}}^{\ell}A\|_{C^{0}}\!\!\leq\!\!(1\!-\!{\hat{C}}\|\chi\|^{-2\tau})\|A\|_{C^{0}}$
holds independently of the value of $i$. Iterating this estimate $\lceil N/\ell\rceil$ times gives the result. ∎
Now we put together the estimates on the behavior of $\mathcal{L}$} on each $H_{\chi}$ and estimate how $\mathcal{L}$} acts on $C^{r}.$
Proof of Theorem 3.5.
First, we decompose $A$ as $\sum_{\chi\in\widehat{M}}A_{\chi}$ where $A_{\chi}\in H_{\chi}.$ Since $A$ has mean 0, $A_{\chi}=0$ when $\chi$ is the trivial character. Thus,
$\|{\mathcal{L}}^{N}A\|_{C^{0}}\leq\sum_{\|\chi\|^{2\tau}<\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}+\sum_{\|\chi\|^{2\tau}\geq\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}.$
For the first term, we will use the spectral gap of $\mathcal{L}$} from Lemma 3.6. This gives us that when $\|\chi\|^{2\tau}<\sqrt{N}$,
$\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}\leq C_{1}\left(1-\frac{C_{2}}{\sqrt{N}}\right)^{N/\ell}\|A\|_{C^{0}}.$
Thus, since there are only $N^{{\dim(M)/4\tau}}$ characters satisfying $\|\chi\|^{2\tau}<\sqrt{N}$,
$\sum_{\|\chi\|^{2\tau}<\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}\leq C_{1}{N^{\dim(M)/4\tau}}\left(1-\frac{C_{2}}{\sqrt{N}}\right)^{N/\ell}\|A\|_{C^{0}}.$
Setting $C_{3}=C_{2}/\ell$ and $a=\dim(M)/4\tau$, this term is $O(N^{a}e^{-C_{3}\sqrt{N}})\ll N^{{(r-\dim(M))}/{4\tau}}$, so up to changing the constant $K$, we have dealt with the first term. Finally, we observe that since $A\in C^{r}$, the projection $A_{\chi}$ satisfies
$\|A_{\chi}\|_{C^{0}}\leq\|A\|_{C^{r}}\|\chi\|^{-r}.$
Therefore, since $\widehat{M}$ corresponds to a lattice in ${\mathbb{R}}^{\dim(M)}$, we can use the classical bounds on tails of higher dimensional $p$-series to see that for some constant $C_{4}>0$,
$\sum_{\|\chi\|^{2\tau}\geq\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}\leq\|A\|_{C^{r}}\sum_{\|\chi\|>N^{1/4\tau}}\frac{1}{\|\chi\|^{r}}\leq C_{4}N^{-(r-\dim(M))/4\tau}\|A\|_{C^{r}}.$
Thus, for the appropriate choice of $K$, $a$, and $b$, the estimate holds. ∎
Theorem 3.7.
Let $G/\Gamma$ be an $m$-great nilmanifold. The set of $m$-tuples in $G^{m}$ which are rapid mixing has full measure with respect to $\mu^{m}$.
Proof.
Whenever an $m$-tuple acts Diophantinely on every level, Theorem 3.5 applies. By (3.3), this is enough to conclude rapid mixing for the random walk they generate. Lemma 3.4, implies that the set of such tuples has full measure with respect to $\mu^{m}$. ∎
4. Nice Words
In this section, we prove Lemma 3.4. To this effect, given $p\geq 1$, let $K_{p}\subset({\mathbb{Z}}^{m})^{(p+1)}$ be the set of all indices $\bf{k_{1},\ldots,k_{p}}$ such that $\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}$ can be written of the form $\log(W_{1}W_{2}^{-1})\text{ mod }\mathfrak{g}^{(p+1)}$ for two words $W_{1}$ and $W_{2}$ of equal length in $\langle F\rangle$.
Lemma 4.1.
$K_{p}$ is Zariski dense.
Proof.
Given an $m$-tuple $F$, choose elements $g_{1},g_{2}\in\langle F\rangle$ and we define new elements $\hat{g}^{(p)}_{L}$ and $\hat{g}^{(p)}_{R}$ inductively as follows. Let $w_{i},w_{i}^{\prime}\in\langle F\rangle$ for each $i\in{\mathbb{N}}_{0}$
(4.1) $\hat{g}_{L}^{(0)}:=w_{0},\quad\hat{g}_{R}^{(0)}:=w_{0}^{\prime},$
(4.2) $\hat{g}_{L}^{(p)}:=\hat{g}_{L}^{(p-1)}w_{p}\hat{g}_{R}^{(p-1)},\quad\hat{g}_{R}^{(p)}:=\hat{g}_{R}^{(p-1)}w_{p}\hat{g}_{L}^{(p-1)}.$
As an initial case, $g_{L}^{(0)},g_{R}^{(0)}]=g_{L}^{(1)}(g_{R}^{(1)})^{-1}.$ By direct computation, we obtain
(4.3) $\begin{split}\hat{g}_{L}^{(p)}(\hat{g}_{R}^{(p)})^{-1}&=\hat{g}_{L}^{(p-1)}w_{p}\hat{g}_{R}^{(p-1)}(\hat{g}_{R}^{(p-1)}w_{p}\hat{g}_{L}^{(p-1)})^{-1}\\ &=[\hat{g}_{L}^{(p-1)}(\hat{g}_{R}^{(p-1)})^{-1},\hat{g}_{R}^{(p-1)}w_{p}]\\ &=[\ldots[[\hat{g}_{L}^{(1)}(\hat{g}_{R}^{(1)})^{-1},\hat{g}_{R}^{(1)}w_{2}],\hat{g}_{R}^{(2)}w_{3}],\ldots],\hat{g}_{R}^{(p-1)}w_{p}]\\ &=[\ldots[[[\hat{g}_{L}^{(0)},\hat{g}_{R}^{(0)}],\hat{g}_{R}^{(1)}w_{2}],\ldots],\hat{g}_{R}^{(p-1)}w_{p}].\end{split}$
By the BCH formula, we write
(4.4) $\log\hat{g}_{L}^{(p)}(\hat{g}_{R}^{(p)})^{-1}=[\ldots[[\log[\hat{g}_{L}^{(0)},\hat{g}_{R}^{(0)}],\log(\hat{g}_{R}^{(1)}w_{2})],\ldots],\log(\hat{g}_{R}^{(p-1)}w_{p})]\text{ mod }\mathfrak{g}^{(p+1)}.$
Letting $\bf{k_{i}}$ be the tuple storing in position $j$ the number of times $g_{j}$ appears in $g_{R}^{(i-1)}w_{i}$, we get that this reduces to
(4.5) $\log\hat{g}_{L}^{(p)}(\hat{g}_{R}^{(p)})^{-1}=[\ldots[\mathbf{k_{0}V,k_{1}V],\ldots,k_{p}V]}\text{ mod }\mathfrak{g}^{(p+1)}.$
Now we wish to understand which sequences $\bf{k_{0},\ldots,k_{p}}$ are possible to choose. Note that $\bf{k_{0}}$ and $\bf{k_{1}}$ are completely free to choose. Meanwhile, for $i\geq 2$, we have $\bf{k_{i}\geq k_{i-1}}$ as the only restraint arising from this construction. It is clear that the set of all tuples satisfying this restraint when interpreted as a subset of ${\mathbb{R}}^{mp}$ is Zariski dense. It follows that $K_{p}$ is a Zariski dense set. ∎
Proof of Lemma 3.4.
We claim that there are words $W_{1}$ and $W_{2}$ such that the coordinates of $\pi_{p}(\log(W_{1}W_{2}^{-1})$ are linearly independent polynomials in $\alpha$. Assume by way of contradiction that for all words $W_{1},W_{2}$, these polynomials are linearly dependent. In this case $K_{p}\subset\mathcal{D}(\mathcal{H}_{m,p}).$ By Lemma 4.1, $\mathcal{H}_{m,p}$ is degenerate contradicting the assumption that $\mathfrak{g}$ is mm–great. The claim follows.
Next, by Theorem 2.5, since the coordinates of $\pi_{p}(W_{1}W_{2}^{-1})$ are linearly independent, for almost every value of the $g_{i}$, $\pi_{p}(W_{1}W_{2}^{-1})\in DC(\gamma,\tau).$ This implies the Lemma. ∎
5. Great Groups
Now that we have established that $m$-greatness implies rapid mixing of almost all random walks on $m$ elements, we wish to show that this condition is not overly restrictive. To that effect, we show that many special Lie algebras are 2-great, and that every step-$s$ Lie algebra is $s$-great. This will prove Theorems 1.2 and 1.3. Finally, we present an example of a group that is not 2-great.
5.1. 2-great algebras
Here we prove that all quasi-abelian, triangular and step-3 or lower Lie algebras are 2-great. Our approach in the first two cases will be to choose a value of $k$ at which $\mathcal{H}_{2,p}(k,\alpha)$ has linearly independent polynomials. We shall use $\bar{k}_{p}=((1,0),(0,1)^{p})$. We will then choose a particular monomial and, by ignoring all variables not appearing in that monomial, find that monomial appears in exactly one coordinate. This will be sufficient to establish linear independence.
Proposition 5.1.
All quasi-abelian groups are 2-great
Proof.
To show this, we must show that $\mathcal{H}_{2,p}$ is nondegenerate for all $1\leq p\leq s$. To that effect, we will let
$V_{1}=\alpha_{0}X+\sum_{j}\alpha_{0,j}Y_{0,j}\mod\mathfrak{g}^{(1)}\quad\text{and}\quad V_{2}=\beta_{0}X+\sum_{j}\beta_{0,j}Y_{0,j}\mod\mathfrak{g}^{(1)}$
and compute the coefficient of $\alpha_{0,i}\beta_{0}^{p}$ in $\mathcal{H}_{2,p}(\bar{k},\alpha).$ In particular, we see that if $(i,p)\in J$ then this coefficient is equal to $Y_{i,p}$. This implies that the polynomials of $\mathcal{H}_{2,p}(\bar{k},\alpha)$ are linearly independent since each one includes a unique monomial. Thus $\mathcal{H}_{2,p}$ is not degenerate for any $p$, and hence any quasi-abelian Lie algebra is 2-great. ∎
Proposition 5.2.
For any $s\in{\mathbb{N}}$, ${\mathbf{t}}_{s}$ is 2-great.
Proof.
Let $V_{1}=\sum_{i=1}^{n-1}\alpha_{0i}E_{i,i+1}\quad\mathrm{mod}\quad\mathfrak{g}^{(1)}$ and $V_{2}=\sum_{i=1}^{n-1}\beta_{0i}E_{i,i+1}\quad\mathrm{mod}\quad\mathfrak{g}^{(1)}.$ We compute the coefficient of $\alpha_{i}\beta_{i+1},\ldots\beta_{i+p}$ in $\mathcal{H}_{2,p}(\bar{k}_{p},\alpha)$ for $1\leq i\leq s-p$ and $1\leq p\leq s$ to be $E_{p,i}$. This implies that the polynomials of $\mathcal{H}_{2,p}(\bar{k},\alpha)$ are linearly independent since each one includes a unique monomial. Thus, we see that $\mathcal{H}_{2,p}$ is not degenerate for any $p$, and conclude that $\mathfrak{t}_{s}$ is 2-great. ∎
We now turn out attention to lower step algebras. We treat first the case of step 1 and step 2 algebras as they may be dealt with quite simply.
Lemma 5.3.
Any abelian or step 2 Lie algebra is 2-great
Proof.
We compute $\mathcal{H}_{2,0}$ and $\mathcal{H}_{2,1}$ for arbitrary groups. If $V_{1}=\sum_{i}\alpha_{i}X_{i}$, $V_{2}=\sum_{i}\beta_{i}X_{i}$, we have that $\mathcal{H}_{2,0}((1,0),\alpha)=V_{1}$ which clearly has linearly independent polynomials for coordinates. Thus $\mathcal{H}_{2,0}$ is always non-degenerate. Now, if $\mathfrak{g}$ is non-abelian, let $X^{(2)}_{j}=[X_{i},X_{i}^{\prime}]$. In $\mathcal{H}_{2,1}$, the monomial $\alpha_{i}\beta_{j}$ only appears as a coefficient of $X^{(2)}_{j}$. Thus, for any non-abelian Lie algebra $\mathcal{H}_{2,1}$ is non-degenerate. From this we conclude the Lemma. ∎
Now, we show that step-3 Lie algebras are also necessarily 2-great. This is the best we can do in terms of step as there exist step-4 Lie algebras that are not 2-great as we shall see.
Proposition 5.4.
Any nilpotent Lie algebra is of step 3 or less is 2-great.
Proof.
We set $V_{1}=\sum_{i}\alpha_{i}X_{i}$, $V_{2}=\sum_{i}\beta_{i}X_{i}$ and show that $M_{121}$ consists of linearly independent polynomials. By way of contradiction, assume that the polynomials of $M_{121}$ are not independent, i.e.
$0\equiv\lambda([[V_{1},V_{2}],V_{1}])\text{ for some }0\neq\lambda\in\mathfrak{g}_{2}^{*}.$
Expanding out, we see that this would imply that
$0\equiv\sum_{i_{1},i_{2},i_{3}\in I}\alpha_{i_{1}}\beta_{i_{2}}\alpha_{i_{3}}\lambda([[X_{i_{1}},X_{i_{2}}],X_{i_{3}}]).$
Since terms of the form $\alpha_{i_{1}}^{2}\beta_{i_{2}}$ occur with coefficient $X_{i_{1}},X_{i_{2}}],X_{i_{1}}]$, we get
(5.1) $\lambda([[X_{i_{1}},X_{i_{2}}],X_{i_{1}}])=0.$
Next expanding the equation $\lambda([[X_{i_{1}}+X_{i_{3}},X_{i_{2}}],X_{i_{1}}+X_{i_{3}}])=0$ as sum of monomials and using (5.1) we get
(5.2) $\lambda([[X_{i_{1}},X_{i_{2}}],X_{i_{3}}]+[[X_{i_{3}},X_{i_{2}}],X_{i_{1}}])=0.$
Combining (5.2) with Jacobi identity
$X_{i_{1}},X_{i_{2}}],X_{i_{3}}]+[[X_{i_{2}},X_{i_{3}}],X_{i_{1}}]+[[X_{i_{3}},X_{i_{1}}],X_{i_{2}}]=0$
we get
(5.3) $\lambda(2[[X_{i_{1}},X_{i_{2}}],X_{i_{3}}]+[[X_{i_{3}},X_{i_{1}}],X_{i_{2}}])=0.$
Swapping $i_{2}$} and $i_{3}$ we obtain
(5.4) $\lambda(2[[X_{i_{1}},X_{i_{3}}],X_{i_{2}}]+[[X_{i_{2}},X_{i_{1}}],X_{i_{3}}])=0.$
Since the matrix defining (5.3)–(5.4) is non degenerate we conclude that
$\lambda([[X_{i_{1}},X_{i_{2}}],X_{i_{3}}])=\lambda([[X_{i_{3}},X_{i_{1}}],X_{i_{2}}])=0.$
Since the equation above holds for arbitrary $i_{1},i_{2},i_{3}\in I$, this implies that $\lambda\!\!=\!\!0$, a contradiction. Thus, $\mathcal{H}_{2,2}$ is non-degenerate. The fact that $\mathcal{H}_{2,0}$ and $\mathcal{H}_{2,1}$ are non-degenerate follows from Lemma 5.3, so any step-3 (or lower) Lie algebra is 2-great. ∎
Combining Propositions 5.1, 5.2, and 5.5 with Theorem 3.7 gives us Theorem 1.3.
5.2. General Lie algebras
We will now show that every nilpotent Lie algebra is mm–great for sufficiently large $m$. Our approach will largely mirror that of the proofs of Propositions 5.1 and 5.2, but taking advantage of the larger value of $m$.
Proposition 5.5.
A step–$s$ Lie algebra is ss–great.
Proof.
We prove that $\mathcal{H}_{s,s-1}$ is non degenerate. Let $V_{i}=\sum_{j=1}^{n_{0}}\alpha_{ij}X_{j}\text{ mod }\mathfrak{g}^{(1)},$ and consider $\bf{[\ldots[k_{0}V,k_{1}V],\ldots],k_{s-1}V]}$. Let $\bar{k}$ denote the conveniently selected parameters: $\mathbf{k_{0}}=(1,0,\ldots,0),\mathbf{k_{1}}=(0,1,\ldots,0),\ldots,\mathbf{k_{s-1}}=(0,0,\ldots,1)$. We compute that the coefficient of $\alpha_{0i_{0}}\cdots\alpha_{s-1i_{s-1}}$ in $\mathcal{H}_{s,s-1}(\bar{k},\alpha)$ is exactly $\ldots[[X_{i_{0}},X_{i_{1}}],\ldots,X_{i_{s-1}}]$. Since each element of the basis for $\mathfrak{g}^{(s-1)}$ can be written of the form $\ldots[[X_{i_{0}},X_{i_{1}}],\ldots,X_{i_{s-1}}]$, we get that each coordinate polynomial of $\mathcal{H}_{s,s-1}(\bar{k},\alpha)$ has a monomial appear which is unique to that coordinate. Thus, the polynomials at $\bar{k}$ are linearly independent, so $\mathcal{H}_{s,s-1}$ is non-degenerate. Since every $\mathfrak{g}/\mathfrak{g}^{(i+1)}$ is step–$s$ or lower, the same argument will show that $\mathcal{H}_{s,i}$ is non degenerate for all $i<s.$ We conclude that $\mathfrak{g}$ is ss–great. ∎
Combining Proposition 5.5 with Theorem 3.7 gives us Theorem 1.2. We also show that $m$-greatness is preserved by taking products and factors.
Proposition 5.6.
The property of $m$-greatness is closed under taking direct products or factors in the following sense:
  • If $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}$ are $m$-great, then $\mathfrak{g}_{1}\times\mathfrak{g}_{2}$ is $m$-great.
  • If $\mathfrak{g}$ is $m$-great and $\mathfrak{h}$ is a factor of $\mathfrak{g}$, then $\mathfrak{h}$ is $m$-great.
Proof.
For clarity, in this proof we will let $\mathcal{H}^{\mathfrak{g}}_{m,p}$ be the map $\mathcal{H}_{m,p}$ associated to the Lie group $\mathfrak{g}.$ Similarly, let $\alpha^{\mathfrak{g}}$ be the Malcev coordinates in the basis of $\mathfrak{g}$ and $\mathcal{D}^{\mathfrak{g}}$ be the set of $k$ such that $\mathcal{H}^{\mathfrak{g}}_{m,p}$ is linearly dependent.
For the first part, observe that $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}})=(\mathcal{H}_{m,p}^{\mathfrak{g}_{1}}(k,\alpha^{\mathfrak{g}_{1}}),\mathcal{H}_{m,p}^{\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{2}}))$. Since both of the $\mathfrak{g}_{i}$’s are $m$-great, $\mathcal{D}_{\mathfrak{g}_{i}}$ is a positive codimension variety. Thus, there exists a $k$ for which $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}}(k,\alpha^{\mathfrak{g}_{1}})$ and $\mathcal{H}_{m,p}^{\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{2}})$ both have linearly independent coordinate polynomials. It follows that for this choice of $k$, $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}})$ has the same property. Thus $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}$ is non-degenerate for any $p$, so $\mathfrak{g}_{1}\times\mathfrak{g}_{2}$ is $m$-great.
For the second part, let $\pi:\mathfrak{g}\rightarrow\mathfrak{h}$ be the factor map. It is straightforward to compute that $\mathcal{H}_{m,p}^{\mathfrak{h}}=\pi\circ\mathcal{H}_{m,p}^{\mathfrak{g}}.$ Thus, the coordinates of $\mathcal{H}^{\mathfrak{h}}_{m,p}$ are the image of the coordinates of $\mathcal{H}^{\mathfrak{g}}_{m,p}$ under an epimorphism. We conclude that since $\mathcal{H}^{\mathfrak{g}}_{m,p}$ was non-degenerate, so is $\mathcal{H}^{\mathfrak{h}}_{m,p}$. Therefore, $\mathfrak{h}$ is $m$-great. ∎
5.3. Counter Example
Although $m$-greatness is a useful property, it has some weaknesses. In particular, although it is expected that on any nilmanifold almost any random walk generated by two translations will be rapid mixing, not all Lie algebras are 22–great. We provide an explicit example of such a Lie algebra.
Example 5.7.
[…]
(5.5) $\begin{split}[X_{1},X_{2}]=Y_{1},\quad[X_{1},X_{3}]=Y_{2},\quad[X_{2},X_{3}]=Y_{3},\\ [Y_{1},X_{1}]=Z_{1},\quad[Y_{1},X_{2}]=Z_{2},\quad[Y_{1},X_{3}]=Z_{3},\\ [Y_{2},X_{1}]=Z_{4},\quad[Y_{2},X_{2}]=Z_{5},\quad[Y_{2},X_{3}]=Z_{6},\\ [Y_{3},X_{1}]=Z_{5}-Z_{3},\quad[Y_{3},X_{2}]=Z_{7},\quad[Y_{3},X_{3}]=Z_{8},\\ [Z_{1},X_{3}]=3W,\quad[Z_{2},X_{3}]=-3W,\quad[Z_{3},X_{1}]=-W,\\ [Z_{3},X_{2}]=W,\quad[Z_{3},X_{3}]=-2W,\quad[Z_{4},X_{2}]=-3W,\\ [Z_{5},X_{1}]=W,\quad[Z_{5},X_{2}]=2W,\quad[Z_{5},X_{3}]=-W,\\ [Z_{6},X_{2}]=3W,\quad[Z_{7},X_{1}]=-3W,\quad[Z_{8},X_{1}]=-3W,\\ [Y_{1},Y_{2}]=4W,\quad[Y_{1},Y_{3}]=-4W,\quad[Y_{2},Y_{3}]=-4W,\end{split}$
while other commutation relations are zero. That these relations actually define a Lie algebra can be verified computationally, but we do not reproduce the calculations here.
Proposition 5.8.
$\mathfrak{g}$ is 2-bad.
Proof.
We compute that $\mathcal{H}_{2,3}\equiv 0$ even though $\mathfrak{g}$ is step 4. By symmetry, it suffices to check that
(5.6) $M_{1211}=[[V_{1},V_{2}],V_{1}],V_{1}]=0,\quad M_{1212}=[[V_{1},V_{2}],V_{1}],V_{2}]=0.$
This may be verified via a computer program, and we present the detailed calculation in Appendix C. ∎
This implies that for any nilmanifold which is a quotient of the Lie group associated with $\mathfrak{g}$, our techniques fail to show rapid mixing of any walk generated by 2 elements. The question of whether such walks are rapid mixing remains open.
Appendix A Lemma for application of Diophantine estimate
The following inequality is useful in estimating the norm of $\mathcal{L}$}.
Lemma A.1.
There exists a $C>0$ such that for all $\theta\in[-1/2,1/2]$,
(A.1) $|1+e^{2\pi i\theta}|\leq 2-C\theta^{2}.$
Proof.
Rearranging, we rewrite (A.1) as
$C\leq\inf_{\theta\in[-1/2,1/2]}\frac{(2-|1+e^{2\pi i\theta}|)}{\theta^{2}}.$
Applying law of cosines and the double angle identity, the inequality becomes
$C\leq\inf_{\theta\in[-1/2,1/2]}\frac{2-2\cos(\pi\theta)}{\theta^{2}}.$
Note that the singularity of the RHS is removable since it can be rewritten as $g(\theta):=2\sum_{n=0}^{\infty}\frac{(\pi\theta)^{2n}}{(2n+2)!}$. Since $g$ is positive and continuous on [1/2,1/2][-1/2,1/2] its achieves its minimum value $C.$
We comment that, in fact, the Lemma holds with $C=8$ but we will not use this in our arguments.
Appendix B Central Limit Theorem
We say that $x_{n}$ satisfies CLT if there is $r>0$ such that for any function $A\in C^{r}(M)$ with zero mean, $\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}A(x_{n})$ converges in distribution to a Gaussian random variable with zero expectations.
Corollary B.1.
If $\mathcal{L}$} satisfies (3.2) then $x_{n}$ satisfies the CLT.
While this result is standard we include the proof for completeness.
Proof.
Let $B=(1-{\mathcal{L}})^{-1}A=\sum_{n=0}^{\infty}{\mathcal{L}}^{n}A.$ By (3.2) this series converges in $C^{0}$ for $r$ large enough. Thus $B(x_{n})=A(x_{n})+{\mathbb{E}}(B(x_{n+1}|{\mathcal{F}}_{n})$ where ${\mathcal{F}}_{n}$ is the $\sigma$-algebra generated by $(x_{0},\dots,x_{n}).$ Summing up, we get $\sum_{n=0}^{N-1}A(x_{n})\!\!=\!\!\sum_{n=1}^{N}\Delta_{n}\!\!+\!\!B(x_{0})\!\!-\!\!B(x_{N})$ where $\Delta_{n}\!\!=\!\!B(x_{n})\!-\!{\mathbb{E}}(B(x_{n})|{\mathcal{F}}_{n-1})$ is a martingale difference sequence. Now by the CLT for martingales (see e.g. [8, §8.2]), to prove the CLT, it is sufficient to show that the limit $\sigma^{2}:=\lim_{N\to\infty}\frac{1}{N}\sum_{n=0}^{N-1}q_{n}$ exists (in probability) where $q_{n}={\mathbb{E}}(\Delta_{n}^{2}|{\mathcal{F}}_{n-1}).$ Note that in our case $q_{n}=Q(x_{n-1})$ for a continuous function $Q$ so the existence of the limit follows from the ergodicity of our Markov chain. ∎
Appendix C Lie algebra calculation
Here we provide the computations relevant to Proposition 5.8. It suffices to check that $V_{1},V_{2}],V_{1}],V_{1}]=[[[V_{1},V_{2}],V_{1}],V_{2}]=0$ since all other triply nested brackets are either necessarily equal to one of these (up to a sign) or forced to be 0. Assume $V_{1}=\sum_{i=1}^{3}\alpha_{i}X_{i},\;\;V_{2}=\sum_{i=1}^{3}\beta_{i}X_{i}$ and we compute the following.
$[[V_{1},V_{2}],V_{1}]=$ $(\alpha_{1}^{2}\beta_{2}-\alpha_{1}\alpha_{2}\beta_{1})Z_{1}+(\alpha_{1}\alpha_{2}\beta_{2}-\alpha_{2}^{2}\beta_{1})Z_{2}+(\alpha_{1}\alpha_{3}\beta_{2}-\alpha_{2}\alpha_{3}\beta_{1})Z_{3}+$
$(\alpha_{1}^{2}\beta_{3}-\alpha_{1}\alpha_{3}\beta_{1})Z_{4}+(\alpha_{1}\alpha_{2}\beta_{3}-\alpha_{2}\alpha_{3}\beta_{1})Z_{5}+(\alpha_{1}\alpha_{3}\beta_{3}-\alpha_{3}^{2}\beta_{1})Z_{6}+$
$(\alpha_{1}\alpha_{2}\beta_{3}-\alpha_{1}\alpha_{3}\beta_{2})(Z_{5}-Z_{3})+(\alpha_{2}^{2}\beta_{3}-\alpha_{2}\alpha_{3}\beta_{2})Z_{7}+(\alpha_{2}\alpha_{3}\beta_{3}-\alpha_{3}^{2}\beta_{2})Z_{8}.$
$[[[V_{1},$ $V_{2}],V_{1}],V_{2}]=\Big((\alpha_{1}^{2}\beta_{2}\beta_{3}-\alpha_{1}\alpha_{2}\beta_{1}\beta_{3})-3(\alpha_{1}\alpha_{2}\beta_{2}\beta_{3}-\alpha_{2}^{2}\beta_{1}\beta_{3})-(\alpha_{1}\alpha_{3}\beta_{1}\beta_{2}-\alpha_{2}\alpha_{3}\beta_{1}^{2})$
$+(\alpha_{1}\alpha_{3}\beta_{2}^{2}-\alpha_{2}\alpha_{3}\beta_{1}\beta_{2})-3(\alpha_{1}\alpha_{3}\beta_{2}\beta_{3}-\alpha_{2}\alpha_{3}\beta_{1}\beta_{3})-3(\alpha_{1}^{2}\beta_{2}\beta_{3}-\alpha_{1}\alpha_{3}\beta_{1}\beta_{2})$
$+(\alpha_{1}\alpha_{2}\beta_{1}\beta_{3}-\alpha_{2}\alpha_{3}\beta_{1}^{2})+2(\alpha_{1}\alpha_{2}\beta_{2}\beta_{3}-\alpha_{2}\alpha_{3}\beta_{1}\beta_{2})-(\alpha_{1}\alpha_{2}\beta_{3}^{2}-\alpha_{2}\alpha_{3}\beta_{1}\beta_{3})$
$+3(\alpha_{1}\alpha_{3}\beta_{2}\beta_{3}-\alpha_{3}^{2}\beta_{1}\beta_{2})-3(\alpha_{2}^{2}\beta_{1}\beta_{3}-\alpha_{2}\alpha_{3}\beta_{1}\beta_{2})-3(\alpha_{2}\alpha_{3}\beta_{1}\beta_{3}-\alpha_{3}^{2}\beta_{1}\beta_{2})$
$+2(\alpha_{1}\alpha_{2}\beta_{1}\beta_{3}-\alpha_{1}\alpha_{3}\beta_{1}\beta_{2})+(\alpha_{1}\alpha_{2}\beta_{2}\beta_{3}-\alpha_{1}\alpha_{3}\beta_{2}^{2})+(\alpha_{1}\alpha_{2}\beta_{3}^{2}-\alpha_{1}\alpha_{3}\beta_{2}\beta_{3})\Big)W=0.$
$[[[V_{1},$ $V_{2}],V_{1}],V_{1}]=\Big(3(\alpha_{1}^{2}\alpha_{3}\beta_{2}-\alpha_{1}\alpha_{2}\alpha_{3}\beta_{1})-3(\alpha_{1}\alpha_{2}\alpha_{3}\beta_{2}-\alpha_{2}^{2}\alpha_{3}\beta_{1})-(\alpha_{1}^{2}\alpha_{3}\beta_{2}-\alpha_{1}\alpha_{2}\alpha_{3}\beta_{1})$
$+(\alpha_{1}\alpha_{2}\alpha_{3}\beta_{2}-\alpha_{2}^{2}\alpha_{3}\beta_{1})-3(\alpha_{1}\alpha_{3}^{2}\beta_{2}-\alpha_{2}\alpha_{3}^{2}\beta_{1})-3(\alpha_{1}^{2}\alpha_{2}\beta_{3}-\alpha_{1}\alpha_{2}\alpha_{3}\beta_{1})$
$+(\alpha_{1}^{2}\alpha_{2}\beta_{3}-\alpha_{1}\alpha_{2}\alpha_{3}\beta_{1})+2(\alpha_{1}\alpha_{2}^{2}\beta_{3}-\alpha_{2}^{2}\alpha_{3}\beta_{1})-(\alpha_{1}\alpha_{2}\alpha_{3}\beta_{3}-\alpha_{2}\alpha_{3}^{2}\beta_{1})$
$+3(\alpha_{1}\alpha_{2}\alpha_{3}\beta_{3}-\alpha_{2}\alpha_{3}^{2}\beta_{1})-3(\alpha_{1}\alpha_{2}^{2}\beta_{3}-\alpha_{1}\alpha_{2}\alpha_{3}\beta_{2})-3(\alpha_{1}\alpha_{2}\alpha_{3}\beta_{3}-\alpha_{1}\alpha_{3}^{2}\beta_{2})$
$+2(\alpha_{1}^{2}\alpha_{2}\beta_{3}-\alpha_{1}^{2}\alpha_{3}\beta_{2})+(\alpha_{1}\alpha_{2}^{2}\beta_{3}-\alpha_{1}\alpha_{2}\alpha_{3}\beta_{2})+(\alpha_{1}\alpha_{2}\alpha_{3}\beta_{3}-\alpha_{1}\alpha_{3}^{2}\beta_{2})\Big)W=0.$
In other words $M_{1211}=M_{1212}=0$, so $\mathcal{H}_{2,3}\equiv 0.$